OpenStudy (highschoolmom2010):

A cube has a net with area 216 in^2. How long is an edge of the cube?

OpenStudy (highschoolmom2010):

@ganeshie8 @gorv

OpenStudy (highschoolmom2010):

i have no idea how to do this

ganeshie8 (ganeshie8):

hint : surface area of cube = 6s^2

OpenStudy (highschoolmom2010):

ok so where do i go with this

ganeshie8 (ganeshie8):

6s^2 = 216 solve s

OpenStudy (highschoolmom2010):

6

OpenStudy (highschoolmom2010):

216/6=6s^2/6 36=s^2 \[\sqrt36=\sqrt s^2\] 6=s

OpenStudy (highschoolmom2010):

like that

ganeshie8 (ganeshie8):

Yup! good job !!

OpenStudy (highschoolmom2010):

thanks for the guidance on this one @ganeshie8

ganeshie8 (ganeshie8):

np :)